# Broad-Line AGN at : The Black Hole Mass Function and a Connection with Little Red Dots

### v1

@@ 4.1 Comparison with Previous Samples
We detect seven of these eight sources in our sample (CEERS-397, CEERS-672, CEERS-746, CEERS-1244, CEERS-1665, and CEERS-2782), but fail to recover one object (CEERS-717) and reject another (CEERS-1236)—the two lowest S/N objects in the final Harikane et al. (2023) sample. Upon examining CEERS-717, we find that the Harikane et al. (2023) custom data reduction extended the spectral coverage of the G395M to 5.3 m, beyond the coverage returned by the standard pipeline (5.2 m). As Harikane et al. (2023) report a spectroscopic redshift of (and thus a line center at 5.21 m) for CEERS-717, this line is too close to the end of the spectral range in our reduction to be robustly detected for our sample. Object CEERS-1236 is initially detected in our broad-line search, but we detect its broad component at S/Nb=3.47, below our threshold of S/N. Harikane et al. (2023) detected this component at S/Nb=5.9, but we expect that this difference may be due to our rescaling of the spectral errors by a factor of 1.7 (see §2.2.2).
In addition to these analyses of CEERS data, Kocevski et al. (2024) previously searched the RUBIES data for spectra of LRDs. In their search, they identified 13 H emitting BLAGN: objects RUBIES-EGS-61496, RUBIES-EGS-60935, RUBIES-EGS-49140, RUBIES-EGS-42232, RUBIES-EGS-42046, RUBIES-EGS-37124, RUBIES-EGS-927271, and RUBIES-EGS-926125 in the CEERS field, and objects RUBIES-UDS-50716 , RUBIES-UDS-47509, RUBIES-UDS-44043, RUBIES-UDS-59971, and RUBIES-UDS-63166 in PRIMER-UDS. We detect 10/13 of these objects, and of these 10, we detect eight with S/N (excluding RUBIES-EGS-927271 S/Nb=2.1, RUBIES-UDS-59971 S/Nb=1.7). For the remaining three objects, the 2D spectrum of RUBIES-EGS-61496 is significantly contaminated such that our optimal flux extraction fails, RUBIES-UDS-50716 fails the curve_fit selection—likely due to poor S/N, and RUBIES-UDS-63166 failed to achieve a S/N5 at its line center pixel. As these objects are likely still BLAGN based on the visual identification of broad components in Kocevski et al. (2024), despite failing our selection procedure, we account for BLAGN such as these through our completeness correction (see §6.1).
Finally, as mentioned above in §3.3, Baggen et al. (2024) previously identified RUBIES-EGS-49140. Beyond these previously known objects, we report a new BLAGN in the CEERS dataset: CEERS-11728, and we report 35 new BLAGN in RUBIES (including the confirmation of a PRISM spectrum candidate BLAGN from Napolitano et al. 2024: RUBIES-EGS-28812).
@@ 5.1 UV and Optical Slopes
To perform these fits, we once again use emcee with a simple power law model: where is the spectral slope, and is an arbitrary normalization constant (using 32 walkers and a 5000 step chain, of which we discard the first 3333 steps). In these initial fits, we do not account for any effects of strong emission lines on the broadband photometry (we address this below and in Figure 8). Using these slopes and our photometry, we apply the full LRD criteria (including , , and half-light radius cuts, as well as color cuts to reject objects with broadband magnitudes significantly effected by emission lines) from Kocevski et al. (2024) to determine that our BLAGN sample contains 10 LRDs. We plot the measured optical slopes () verses the measured UV slopes () of both the LRD and non-LRD BLAGN populations in Figure 7.
We also find nine objects that satisfy the LRD and criteria, but do not meet the full Kocevski et al. (2024) LRD selection. Eight of these objects are classified as non-LRDs due to cuts designed to remove objects with broadband colors that might be significantly effected by strong emission lines (see Kocevski et al. 2024, §3.1 for details). The remaining object is insufficiently compact in F444W imaging. The overlap of our full BLAGN sample with the LRD population is small (20%), indicating clearly that LRDs do not account for all BLAGN at . These non-LRD BLAGN objects show diverse UV and optical slopes, likely indicating a variety of reddenings (or lack thereof) of both AGN and nebular emission. For a detailed analysis of reddening of the individual objects in this sample, see Brooks et al. (2024; in prep).
@@
Figure 8 shows that while the six BLAGN LRDs for which we have PRISM spectra all show signs of emission line boosting, four out of six still exhibit red . This implies that a significant fraction of the LRDs are intrinsically dust-reddened in the optical beyond the contributions of emission lines to their broadband colors. Interestingly, Figure 8 also shows that seven objects rejected by Kocevski et al. (2024) due to possible emission line contamination actually exhibit continua with . As such, photometric definitions and criteria for LRD selection may be both slightly contaminated (as seen by the two out of seven LRDs that have spectroscopic ) and moderately incomplete (as seen by the seven photometric non-LRDs that have spectroscopic ). For simplicity, and to best comment on, analyze, and compare to the populations of LRDs identified in photometric searches, we continue to use the photometric Kocevski et al. (2024) criteria throughout this study, and reserve a dedicated investigation into the properties of spectroscopic LRDs for a future work.
@@ 5.2 Stacked H Lines
We next produce stacks of both the 10 LRD and 40 non-LRD BLAGN H lines to better compare the broad-line properties of the two populations. Here we shift each object’s G395M spectrum to the rest frame by dividing both its wavelength and flux values by , where is the spectroscopic redshift measured from the peak of H line. We next resample these rest-frame spectra to a common wavelength grid while preserving the observed frame spectral resolution. To best analyze the shape of the H line without biasing our stacks toward the brightest lines, we next re-normalize each object’s spectrum to the measured (broad+narrow component) peak of the H line . Finally, we combine the re-normalized spectra with a simple, unweighted, resampled pixel-by-pixel mean (and combine the associated errors in quadrature). We show these stacked H lines in Figure 9
We next fit these stacked lines using the same two-component H (with [Nii]) model as in the single object fitting. We summarize the results of the fits in Table 2. These measurements show that the LRDs tend to have broad components that are both broader ( versus km s-1) and make up a larger fraction of the total H line emission (75% vs 58%) than the broad components of the non-LRDs. Moreover, the LRDs show little to no significant [Nii] emission, while the non-LRDs exhibit a modest log10([Nii]6585 / H) of . These measurements imply that observations of H in the LRD-selected BLAGN are more significantly dominated by emission from the broad-line region, with lesser contributions from narrow emission than non-LRD BLAGN.
@@ 6 The Black Hole Mass Function
We further verify our black hole masses by re-deriving them using the prescription from Greene & Ho (2005). Here, we combine their equations (3) and (5) to derive an expression (our equation 2) for black hole mass as a function of continuum luminosity at rest-frame 5100 Å () and the FWHM of the broad component of H (). We measure directly from the PRISM spectra (when available) and use our fitted values of FWHM in this expression. We find that the derived black hole masses are well-matched to the black hole masses with a median offset and an absolute median deviation of 0.22. As this absolute median deviation is still significantly less than the assumed dex uncertainty inherent in these prescriptions, we consider this to be good agreement. With this confirmation, and our lack of PRISM data for some objects, we use our black hole masses in the remainder of this work.
@@
We next estimate the effects of dust reddening on our computed BH masses using the methodology from Kocevski et al. (2024) (see their §3.3 for full details). Briefly, we use a custom fitter to fit composite AGN and star-formation models to the photometric data for each source. From these fits, we derive an approximate reddening for each AGN. We caution that these fits may be significantly skewed by the presence of strong emission lines and assume that the AGN component dominates the reddened rest-frame optical continuum emission. As such, we use these fits as simply estimates of . We next assume a Calzetti et al. (2000) dust law (with ), correct our H fluxes for this reddening, and recompute for each object. Based on these estimates, we find that the median increase in for the BLAGN sample is 0.21 dex, with a maximum increase of 0.54 dex for the most reddened object in the dataset. This effect is most prominent for the more massive objects in the dataset, where objects with have a median correction of 0.05 dex, and objects with have a median correction of 0.32 dex. While these corrections offer an estimate of the un-attenuated H fluxes, due to the uncertainty in our SED modelling and fitting, and the complexity of the SEDs of LRDs and BLAGN as a whole, we again continue to use our uncorrected BH masses and H fluxes in the remainder of this work with the caveat that the BH masses of our BLAGN population may be underestimated by dex at and by dex at . For a detailed analysis of reddening of the individual objects in this sample, see Brooks et al. (2024; in prep).
@@
We plot the computed BH masses as a function of object redshift in Figure 10 (with the 0.5 dex systematic uncertainty depicted as a black error bar) and list the computed masses in Table 1. Notably, the BH masses estimated here are in broad agreement with those found in other collections of JWST-detected BLAGN at (Harikane et al., 2023; Maiolino et al., 2023a) ranging from .
@@
To derive our selection volume, we require a measurement of our survey area. For this, we use the full areas of the four quadrants of the NIRSpec MSA for each pointing to determine the field areas (shown as black rectangles in Figure 6), and after accounting for overlap between NIRSpec pointings, we determine a total probed area of 155.87 arcmin2 for the 21 NIRSpec MSA pointings used in this study (six from CEERS, six from RUBIES-EGS, and nine from RUBIES-UDS). We choose a redshift range of for our BH mass function that encompasses the main distribution of BLAGN in our sample (this excludes the two highest redshift sources at and ). The co-moving volume over this redshift range and area is 1.37 106 Mpc3.
@@ 6.1 Sample Completeness
To quantify the prevalence of these missed sources, we use a simulation using the Pandeia JWST ETC engine (Pontoppidan et al., 2016) to produce simulated JWST NIRSpec MSA G395M spectra featuring modeled broad H lines. These modelled H lines use a seven parameter model consisting of: narrow H flux, broad H flux, narrow H FWHM, broad H FWHM, [Nii]/H ratio, H equivalent width (EW, computed with the sum of the broad and narrow components, and used to set the continuum flux level), and redshift. Critically, this combination of parameters also inherently includes a derived BH mass for each synthetic source. We vary these parameters on a large seven-dimensional grid and produce a sample of broad H lines that span the physical parameters of our BLAGN sample. We sample narrow H FWHM from 1–750 km s-1, broad H FWHM from 500–3000 km s-1, [Nii]/H ratio from 0–0.5, and redshift from 3.5–6.0 as flat distributions in linear space. We sample narrow and broad H flux from ergs s-1 cm-2 and H rest-frame equivalent width from 100–1000 Å as flat distributions in logarithmic space. While this grid may not be representative of the true underlying distribution of BLAGN, as we do not have any strong priors on the H line properties of BLAGN at these redshifts and line fluxes outside of our own detected sample, we use the above simple flat priors grid as our best option, and simulate objects.
We adopt the RUBIES NIRSpec observation strategy and parameters and use the modelled H spectra as input spectra for Pandeia to produce simulated observed spectra. We then analyze this sample of synthetic spectra using the same broad-line detection/fitting algorithms and parameters as described in §3.1.
@@
We find that the dominant variables in determining an object’s detection are the flux of the narrow and broad H components, while the redshift of the source, the [N ii]/H ratio and the H EW have minimal to no effect. This is fully expected, as in order to be initially detected, an object must have a peak pixel S/N, which is dependant on both the broad and narrow component fluxes. Then, the object must ultimately fit with a broad component S/N, which depends directly on the broad component flux. These effects are also coupled to the FWHM of the broad and narrow components, such that broader lines will be harder to detect, as the flux is spread over larger areas of the spectrum and thus incur a greater amount of read-noise and a lower S/N on the line peak. This effect explains why the completeness trends steadily downwards with decreasing instead of dropping suddenly from near 100% to near 0% at a cutoff corresponding to a strict S/N cut.
@@ 6.3 Comparison with Empirical and Theoretical Models
Here, we briefly summarize the key aspects of this modeling that are most relevant for this work, and refer the reader to the original studies for details. First, we show results from two semi-analytic models of the co-evolution of galaxies and SMBH, which are based within dark matter merger histories. The Santa Cruz semi-analytic model (SC-SAM; Somerville et al., 2008; Hirschmann et al., 2012; Yung et al., 2021) populates top level halos with seeds of mass , where top level halos are 0.01 times the mass of the root halo or , whichever is smaller. Rapid black hole growth is triggered by mergers and disk instabilities, where the growth rate is capped at the Eddington rate. Following a triggering event, BHs grow at the Eddington rate until they reach a critical mass, above which BH feedback halts further growth. BHs can also grow by merging. The Trinca et al. (2022) semi-analytic model (Cosmic Archaeology Tool; CAT) explored different combinations of BH seeding and accretion scenarios, including both light (Pop III) and heavy (direct collapse; ) seeds, and both Eddington-limited accretion and a model in which brief periods of super-Eddington accretion occur following galaxy mergers. We also show results from two semi-empirical/analytic models. Li et al. (2024a) seed BHs at and apply a parameterized stochastic accretion model. They calibrated the parameters controlling seeding and accretion to reproduce the observed quasar luminosity function at lower redshifts. We further consider the distribution of BH masses derived from a Press-Schechter halo mass function, employing a conversion efficiency between halo and BH masses, which in turn assumes that the BH feedback energy is balanced by the gravitational potential energy of the host system (Jeon et al., 2022). Lastly, we show the BH mass function from the FLARES cosmological hydrodynamical simulation (Wilkins et al., 2024). BH seeds with a mass of h are inserted into halos with masses greater than , and accretion is modeled with an Eddington-limited Bondi-Hoyle approach. BHs can also grow by merging. Thermal energy associated with black hole accretion is injected stochastically into particles surrounding the BH.
@@
All of the above mentioned models and simulations predict a rather featureless, power law BH mass function over the observed range of BH masses. Even more strikingly, the three most detailed models (the SC-SAM, CAT, and FLARES) reproduce the observed BH mass function shape and normalization quite well over this mass range, and the other models are also in the right general ballpark. This is significant because many models have difficulty growing large enough populations of the more massive black holes at (with masses of a few billion ) from physically motivated seed masses with Eddington-limited accretion. This comparison shows that, although the newly discovered population of lower mass BHs uncovered by JWST is often described as surprising or unexpected, it is actually fully consistent with theoretical expectations from several models (see also: Ricarte & Natarajan, 2018).

### v2

@@ 4.1 Comparison with Previous Samples
We detect seven of these eight sources in our sample (CEERS-397, CEERS-672, CEERS-746, CEERS-1244, CEERS-1665, and CEERS-2782), but fail to recover one object (CEERS-717) and reject two others (CEERS-1236, CEERS-1665). Upon examining CEERS-717, we find that the Harikane et al. (2023) custom data reduction extended the spectral coverage of the G395M to 5.3 m, beyond the coverage returned by the standard pipeline (5.2 m). As Harikane et al. (2023) report a spectroscopic redshift of (and thus a line center at 5.21 m) for CEERS-717, this line is too close to the end of the spectral range in our reduction to be robustly detected for our sample. Object CEERS-1236 is initially detected in our broad-line search, but we detect its broad component at S/Nb=3.47, below our threshold of S/N. Harikane et al. (2023) detected this component at S/Nb=5.9, but we expect that this difference may be due to our rescaling of the spectral errors by a factor of 1.7 (see §2.2.2). We also initially detect Object CEERS-1665 as a broad-line candidate, but like CEERS-1236, we detect its broad component at only S/Nb=3.86: below our S/N cutoff. Harikane et al. (2023) detected this broad component with S/Nb=11.6 and FWHMb=1794 km s-1, but our fit detects the weak broad component with FWHMb=838 km s-1. In a visual inspection of the two fitted spectra, both fits seem to match their respective data well, therefore we attribute this discrepancy to differences in data reduction or continuum modeling (or lack thereof) during the line fitting.
In addition to these analyses of CEERS data, Kocevski et al. (2024) previously searched the RUBIES data for spectra of LRDs. In their search, they identified 13 H emitting BLAGN: objects RUBIES-EGS-61496, RUBIES-EGS-60935, RUBIES-EGS-49140, RUBIES-EGS-42232, RUBIES-EGS-42046, RUBIES-EGS-37124, RUBIES-EGS-927271, and RUBIES-EGS-926125 in the CEERS field, and objects RUBIES-UDS-50716 , RUBIES-UDS-47509, RUBIES-UDS-44043, RUBIES-UDS-59971, and RUBIES-UDS-63166 in PRIMER-UDS. We detect 11/13 of these objects, and of these 11, we detect eight with S/N (excluding RUBIES-EGS-927271 S/Nb=1.9, RUBIES-UDS-59971 S/Nb=3.1, RUBIES-UDS-50716). For the remaining two objects, the 2D spectrum of RUBIES-EGS-61496 is significantly contaminated such that our optimal flux extraction fails and RUBIES-UDS-63166 failed to achieve a S/N5 at its line center pixel. As these objects are likely still BLAGN based on the visual identification of broad components in Kocevski et al. (2024), despite failing our selection procedure, we account for BLAGN such as these through our completeness correction (see §6.1).
Finally, as mentioned above in §3.3, Wang et al. (2024b) previously identified RUBIES-EGS-49140. Beyond these previously known objects, we report a new BLAGN in the CEERS dataset: CEERS-11728, and we report 49 new BLAGN in RUBIES (including the confirmation of a PRISM spectrum candidate BLAGN from Napolitano et al. 2024: RUBIES-EGS-28812).
@@ 5.1 UV and Optical Slopes
To perform these fits, we once again use emcee with a simple power law model: where is the spectral slope, and is an arbitrary normalization constant (using 32 walkers and a 5000 step chain, of which we discard the first 3333 steps). In these initial fits, we do not account for any effects of strong emission lines on the broadband photometry (we address this below and in Figure 8). Using these slopes and our photometry, we apply the full LRD criteria (including , , and half-light radius cuts, as well as color cuts to reject objects with broadband magnitudes significantly effected by emission lines) from Kocevski et al. (2024) to determine that our BLAGN sample contains 21 LRDs. We plot the measured optical slopes () verses the measured UV slopes () of both the LRD and non-LRD BLAGN populations in Figure 7.
We also find three objects that satisfy the LRD and criteria, but do not meet the full Kocevski et al. (2024) LRD selection. All three of these objects are classified as non-LRDs due to cuts designed to remove objects with broadband colors that might be significantly effected by strong emission lines (see Kocevski et al. 2024, §3.1 for details). The remaining object is insufficiently compact in F444W imaging. The overlap of our full BLAGN sample with the LRD population is moderate (34%), indicating clearly that LRDs do not account for all BLAGN at . These non-LRD BLAGN objects show diverse UV and optical slopes, likely indicating a variety of reddenings (or lack thereof) of both AGN and nebular emission. For a detailed analysis of reddening of the individual objects in this sample, see (Brooks et al., 2024).
@@
Figure 8 shows that only two of the BLAGN LRDs for which we have PRISM spectra show strong signs of emission line boosting. This implies that a significant fraction of the LRDs are indeed intrinsically dust-reddened in the optical beyond the contributions of emission lines to their broadband colors. Interestingly, Figure 8 also shows that seven objects rejected by Kocevski et al. (2024) due to possible emission line contamination actually exhibit continua with . As such, photometric definitions and criteria for LRD selection may be both slightly contaminated (as seen by the two out of 16 LRDs that have spectroscopic ) and moderately incomplete (as seen by the seven photometric non-LRDs that have spectroscopic ). For simplicity, and to best comment on, analyze, and compare to the populations of LRDs identified in photometric searches, we continue to use the photometric Kocevski et al. (2024) criteria throughout this study, and reserve a dedicated investigation into the properties of spectroscopic LRDs for a future work.
@@ 5.2 Stacked H Lines
We next produce stacks of both the 21 LRD and 42 non-LRD BLAGN H lines to better compare the broad-line properties of the two populations. Here we shift each object’s G395M spectrum to the rest frame by dividing both its wavelength and flux values by , where is the spectroscopic redshift measured from the peak of H line. We next resample these rest-frame spectra to a common wavelength grid while preserving the observed frame spectral resolution. To best analyze the shape of the H line without biasing our stacks toward the brightest lines, we next re-normalize each object’s spectrum to the measured (broad+narrow component) peak of the H line. Finally, we combine the re-normalized spectra with a simple, unweighted, resampled pixel-by-pixel mean (and combine the associated errors in quadrature). We show these stacked H lines in Figure 9
We next fit these stacked lines using the same two-component H (with [Nii]) model as in the single object fitting. We summarize the results of the fits in Table 2 and show the fits in Figure 9. These measurements show that the LRDs tend to have broad components that are both broader ( versus km s-1) and make up a larger fraction of the total H line emission (73% vs 59%) than the broad components of the non-LRDs. Moreover, the LRDs show little to no significant [Nii] emission, while the non-LRDs exhibit a modest log10([Nii]6585 / H) of . These measurements imply that observations of H in the LRD-selected BLAGN are more significantly dominated by emission from the broad-line region, with lesser contributions from narrow emission than non-LRD BLAGN.
We repeat this stacking process (this time using inverse variance weighting and no normalization) for the [O iii] doublet to retest both the LRD and non-LRD populations for the presence of outflows. We apply the same fitting procedures as described in §3.3, and find that neither stack shows evidence for outflows. Specifically, we measure a BIC=4.9 in favor of a outflow component model for the [Oiii] doublet in the LRD stack with a outflow S/N=0.34. The non-LRD stack exhibits a BIC=1.6 in favor of an outflow component, and the component is also insufficiently detected with a S/N=0.42. As such, we do not find significant evidence for outflows in our LRD or non-LRD samples.
@@ 6 The Black Hole Mass Function
We further verify our black hole masses by re-deriving them using the prescription from Greene & Ho (2005). Here, we combine their equations (3) and (5) to derive an expression (our equation 2) for black hole mass as a function of continuum luminosity at rest-frame 5100 Å () and the FWHM of the broad component of H (). We measure directly from the PRISM spectra (when available) and use our fitted values of FWHM in this expression. We find that the derived black hole masses are well-matched to the black hole masses with a median offset 0.13 and an absolute median deviation of 0.28. As this absolute median deviation is still significantly less than the assumed dex uncertainty inherent in these prescriptions, we consider this to be good agreement. With this confirmation, and our lack of PRISM data for some objects, we use our black hole masses in the remainder of this work.
@@
We next estimate the effects of dust reddening on our computed BH masses using the methodology from Kocevski et al. (2024) (see their §3.3 for full details). Briefly, we use a custom fitter to fit composite AGN and star-formation models to the photometric data for each source. From these fits, we derive an approximate reddening for each AGN. We caution that these fits may be significantly skewed by the presence of strong emission lines and assume that the AGN component dominates the reddened rest-frame optical continuum emission. As such, we use these fits as simple estimates of . We next assume a Calzetti et al. (2000) dust law (with ), correct our H fluxes for this reddening, and recompute for each object. Based on these estimates, we find that the median increase in for the BLAGN sample is 0.19 dex, with a maximum increase of 0.55 dex for the most reddened object in the dataset. This effect is most prominent for the more massive objects in the dataset, where objects with have a median correction of 0.05 dex, and objects with have a median correction of 0.31 dex. While these corrections offer an estimate of the un-attenuated H fluxes, due to the uncertainty in our SED modeling and fitting, and the complexity of the SEDs of LRDs and BLAGN as a whole, we again continue to use our uncorrected BH masses and H fluxes in the remainder of this work with the caveat that the BH masses of our BLAGN population may be underestimated by dex at and by dex at . For a detailed analysis of reddening of the individual objects in this sample, see Brooks et al. (2024). In brief, Brooks et al. (2024) examine this sample, as well as Harikane et al. (2023); Kocevski et al. (2023), and Maiolino et al. (2023a), and find that the broad-line component of H is undetected () in over 86% of JWST identified BLAGN. Therefore, it is very difficult to constrain the reddenings of the broad-line regions of these objects using Balmer lines (see Brooks et al., 2024, their Figure 4). Nonetheless, our analysis above indicates that the effect of reddening on our computed BH masses (+0.19 dex) is small compared to the 0.5 dex uncertainty from the Reines et al. (2013) prescription.
@@
We plot the computed BH masses as a function of object redshift in Figure 10 (with the 0.5 dex systematic uncertainty depicted as a black error bar) and list the computed masses in Table 1. Notably, the BH masses estimated here are in broad agreement with those found in other collections of JWST-detected BLAGN at (e.g., Carnall et al., 2023; Harikane et al., 2023; Maiolino et al., 2023b; Greene et al., 2024; Juodžbalis et al., 2024; Killi et al., 2023; Kocevski et al., 2024; Lin et al., 2024; Übler et al., 2024; Wang et al., 2024a) ranging from . We also note that while our sample of LRDs span roughly the same range of BH masses as the non-LRD sample, as a population, the LRDs exhibit greater BH masses than the non-LRDs . However, as there are only 21 LRDs in our sample, our sample is still too small for a robust statistical comparison of these mass distributions.
@@
To derive our selection volume, we require a measurement of our survey area. For this, we use the full areas of the four quadrants of the NIRSpec MSA for each pointing to determine the field areas (shown as black rectangles in Figure 6), and after accounting for overlap between NIRSpec pointings, we determine a total probed area of 181.98 arcmin2 for the 24 NIRSpec MSA pointings used in this study (six from CEERS, six from RUBIES-EGS, and 12 from RUBIES-UDS). We choose a redshift range of for our BH mass function that encompasses the main distribution of BLAGN in our sample (this excludes the two highest redshift sources at and ). The co-moving volume over this redshift range and area is 1.60 106 Mpc3.
@@ 6.1 Sample Completeness
To quantify the prevalence of these missed sources, we use a simulation using the Pandeia JWST ETC engine (Pontoppidan et al., 2016) to produce simulated JWST NIRSpec MSA G395M spectra featuring modeled broad H lines. These modeled H lines use a seven parameter model consisting of: narrow H flux, broad H flux, narrow H FWHM, broad H FWHM, [Nii]/H ratio, H equivalent width (EW, computed with the sum of the broad and narrow components, and used to set the continuum flux level), and redshift. Critically, this combination of parameters also inherently includes a derived BH mass for each synthetic source. We vary these parameters on a large seven-dimensional grid and produce a sample of broad H lines that span the physical parameters of our BLAGN sample. We sample narrow H FWHM from 1–750 km s-1, broad H FWHM from 500–3000 km s-1, [Nii]/H ratio from 0–0.5, and redshift from 3.5–6.0 as flat distributions in linear space. We sample narrow and broad H flux from ergs s-1 cm-2 and H rest-frame equivalent width from 100–1000 Å as flat distributions in logarithmic space. While this grid may not be representative of the true underlying distribution of BLAGN, as we do not have any strong priors on the H line properties of BLAGN at these redshifts and line fluxes outside of our own detected sample, we use the above simple flat priors grid as our best option, and simulate objects.
We adopt the RUBIES NIRSpec observation strategy and parameters and use the modeled H spectra as input spectra for Pandeia to produce simulated observed spectra. We then analyze this sample of synthetic spectra using the same broad-line detection/fitting algorithms and parameters as described in §3.1.
@@
We find that the dominant variables in determining an object’s detection are the flux of the narrow and broad H components, while the redshift of the source, the [N ii]/H ratio and the H EW have minimal to no effect. This is fully expected, as in order to be initially detected, an object must have a peak pixel S/N, which is dependent on both the broad and narrow component fluxes. Then, the object must ultimately fit with a broad component S/N, which depends directly on the broad component flux. These effects are also coupled to the FWHM of the broad and narrow components, such that broader lines will be harder to detect, as the flux is spread over larger areas of the spectrum and thus incur a greater amount of read-noise and a lower S/N on the line peak. This effect explains why the completeness trends steadily downwards with decreasing instead of dropping suddenly from near 100% to near 0% at a cutoff corresponding to a strict S/N cut.
@@ 6.3 Comparison with Empirical and Theoretical Models
Here, we briefly summarize the key aspects of this modeling that are most relevant for this work, and refer the reader to the original studies for details. First, we show results from two semi-analytic models of the co-evolution of galaxies and SMBH, which are based within dark matter merger histories. The Santa Cruz semi-analytic model (SC-SAM; Somerville et al., 2008; Hirschmann et al., 2012; Yung et al., 2021) populates top level halos with seeds of mass , where top level halos are 0.01 times the mass of the root halo or , whichever is smaller. Rapid black hole growth is triggered by mergers and disk instabilities, where the growth rate is capped at the Eddington rate. Following a triggering event, BHs grow at the Eddington rate until they reach a critical mass, above which BH feedback halts further growth. BHs can also grow by merging. The Cosmic Archaeology Tool (CAT) semi-analytic model (Trinca et al., 2022) explored different combinations of BH seeding and accretion scenarios, including both light (Pop III) and heavy (direct collapse; ) seeds, and both Eddington-limited accretion and a model in which brief periods of super-Eddington accretion occur following galaxy mergers. We also show results from two semi-empirical/analytic models. Li et al. (2024a) seed BHs at and apply a parameterized stochastic accretion model. They calibrated the parameters controlling seeding and accretion to reproduce the observed quasar luminosity function at lower redshifts. We further consider the distribution of BH masses derived from a Press-Schechter halo mass function, employing a conversion efficiency between halo and BH masses, which in turn assumes that the BH feedback energy is balanced by the gravitational potential energy of the host system (Jeon et al., 2022). Additionally, we show the BH mass function from the FLARES cosmological hydrodynamical simulation (Wilkins et al., 2024). BH seeds with a mass of h are inserted into halos with masses greater than , and accretion is modeled with an Eddington-limited Bondi-Hoyle approach. BHs can also grow by merging. Thermal energy associated with black hole accretion is injected stochastically into particles surrounding the BH. Finally, we show the BH mass function from the empirical Trinity model (Zhang et al., 2023). Trinity parametrizes the dark matter halo—galaxy–SMBH connection from z=0-10 as a series of redshift-dependent scaling relations, and finds the best fitting parameters that simultaneously match a large compilation of galaxy data from and SMBH data from z=0-6.5. From the best-fitting Trinity model, we extract the SMBH mass—luminosity distributions to forward model the BLAGN mass functions.
@@
All of the above mentioned models and simulations predict a rather featureless, power law BH mass function over the observed range of BH masses. Even more strikingly, the three most detailed models (the SC-SAM, CAT, and FLARES) reproduce the observed BH mass function shape and normalization quite well over this mass range, and the other models are also in the right general ballpark. This is significant because many models have difficulty growing large enough populations of the more massive black holes at (with masses of a few billion ) from physically motivated seed masses with Eddington-limited accretion. This comparison shows that, although the newly discovered population of lower mass BHs uncovered by JWST is often described as surprising or unexpected, it is actually consistent with theoretical expectations from several models (see also: Ricarte & Natarajan, 2018). However, this consistency comes with the caveat that all of the above BH mass function models show the total BH mass function consisting of all galaxies, not just those hosting BLAGN. As a comparison, we show a modified BH mass function in Figure 13 (orange dotted curve) from Trinity that only shows the contribution of BLAGN to the BH mass function. This BLAGN-only BH mass function shows a 1-2 dex separation from our observed BLAGN BH mass function, suggesting that—especially at the low-mass end—in at least this particular model, the BLAGN mass function may underestimate either the active fraction of SMBHs or the total mass function when compared to JWST-based observations.